# Disclosure of AI-Generated News Increases Engagement but Does Not Reduce Aversion, Despite Positive Quality Ratings

### v2

@@ Abstract
The advancement of artificial intelligence (AI) has led to its application in many areas, including news media. The integration of AI in journalism presents both opportunities and risks for democracy, making it crucial to understand public reception of and engagement with AI-generated news, as it may directly influence political knowledge and trust. This preregistered study investigates (i) the perceived quality of AI-assisted and AI-generated versus human-generated news articles, (ii) whether disclosure of AI’s involvement in generating these news articles influences engagement with them, and (iii) whether such awareness affects the willingness to read AI-generated articles in the future. We employed a between-subjects survey experiment with 599 participants from the German-speaking part of Switzerland, who evaluated the credibility, readability, and expertise of news articles. These articles were either written by journalists (control group), rewritten by AI (AI-assisted group), or entirely generated by AI (AI-generated group). Our results indicate that all news articles, regardless of whether they were written by journalists or AI, were perceived to be of equal quality. When participants in the treatment groups were subsequently made aware of AI’s involvement in generating the articles, they expressed a higher willingness to engage with (i.e., continue reading) the articles than participants in the control group. However, they were not more willing to read AI-generated news in the future. These results suggest that aversion to AI usage in news media is not primarily rooted in a perceived lack of quality, and that by disclosing using AI, journalists could attract more immediate engagement with their content, at least in the short term.
@@ 1 Introduction
The news media landscape is evolving significantly with the advancement of artificial intelligence (AI). AI applications, especially in natural language generation, have become integral to the production of news (Cools and Diakopoulos, 2024; Simon, 2024b, a). The integration of AI in the news presents both opportunities and risks for democracy (Arguedas and Simon, 2024). On the one hand, AI has the potential to make news more accessible and tailored to diverse audiences, supporting informed public debate. On the other hand, however, challenges arise around transparency, bias, and trust, as AI models often reflect the dominant worldviews of their creators, potentially narrowing the range of perspectives in public discourse.
@@
A specific issue is the public’s perception of AI-produced news. Understanding attitudes towards AI-generated news is essential as it may directly impact the public’s political knowledge and trust. Indeed, if the public distrusts AI-generated content and chooses not to engage with it, they may miss out on important political information, leading to a less informed citizenry. A recent study revealed that only 29% of Swiss respondents would read fully AI-generated news, 55% would read news if AI assisted in its creation, and 84% would prefer news created without AI involvement (Vogler et al., 2023). This highlights skepticism towards AI-generated news amongst the Swiss population, consistent with surveys from other countries (Fletcher and Nielsen, 2024). Such aversion could be problematic because it could undermine the credibility and effectiveness of news organizations. Therefore, transparent and responsible use of AI in journalism is essential to build trust and ensure that the public remains informed and engaged with credible news sources.
@@
In this preregistered study (see the Appendix for an anonymized version of the pre-registration), conducted with Swiss participants, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future. Our results indicate that all kinds of news articles, regardless of whether they were written by journalists or (assisted by) AI, were perceived to be of similar quality in terms of credibility, readability, and expertise. When participants in the treatment groups were subsequently informed about AI’s role in generating the articles, they expressed a higher willingness to read the full news article than participants in the control group, who saw human-written articles. This increased engagement, however, did not translate to an increased openness towards reading AI-generated news in the future.
These findings imply that aversion towards AI usage in the news media is not primarily linked to concerns that AI-generated content is perceived to be of lower quality. Moreover, the results suggest that transparency in the use of AI by journalists could enhance immediate reader engagement. Future work should delve deeper into understanding the underlying mechanisms driving both results, exploring factors such as the novelty of AI involvement, trust in AI, and the psychological effects of transparency. Additionally, research should investigate long-term impacts on reader behavior and trust to develop strategies for effectively integrating AI in journalism, ensuring that it complements rather than undermines the journalistic process.
2 Previous Research
@@
AI is increasingly integrated into news production workflows, handling tasks such as data analysis, trend identification, content aggregation, and personalization (Cools and Diakopoulos, 2024). Structured data-based formats like sports and financial reporting are particularly amenable to AI-generated content due to their formulaic nature. AI can assist with data-heavy tasks, enabling journalists to focus on responsibilities that demand human judgment, creativity, and critical thinking, such as investigative reporting and nuanced writing (Clerwall, 2014; Haim and Graefe, 2017). These applications suggest potential for further AI expansion into other news formats, which may amplify the news industry’s dependency on technological platforms (Simon, 2024b).
Research on public perceptions of news authored by AI reveals a mixed picture. While earlier studies demonstrated preferences for human-written news (Graefe et al., 2018; Graefe and Bohlken, 2020), more recent work found few notable differences in perceived quality, expertise, credibility, and readability between human-written and AI-generated content (Wölker and Powell, 2021; Kim et al., 2020; Kim, Xu and Merrill Jr, 2022). However, AI-produced content tend to be distrusted, especially for politically or emotionally sensitive subjects (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020). People tend to view AI-produced news as less trustworthy and transparent, and less worth paying for over human-written news (Fletcher and Nielsen, 2024). Most support labeling AI involvement, especially for core tasks like writing, with more acceptance in soft news topics than in hard news. However, public perceptions of AI-generated news often depend on the specific ways AI is implemented and whether or how its use is disclosed (Altay and Gilardi, 2024). AI can be viewed as more objective and less biased than human journalists (Wölker and Powell, 2021; Wang and Ophir, 2024). However, perceptions of algorithmic neutrality can be misleading, as AI systems trained on biased data may inadvertently perpetuate or amplify existing biases.
@@
While transparency is generally seen as beneficial for accountability and audience trust (Wang and Ophir, 2024), simply disclosing AI’s involvement may be insufficient or even counterproductive (Wang and Huang, 2024; Wang and Ophir, 2024). People tend to be skeptical of headlines labeled as AI-generated, regardless of whether they are genuinely AI-produced or actually created by humans, partly because they tend to assume an extreme, possibly unrealistic degree of AI involvement (Altay and Gilardi, 2024). Without careful management, transparency can lead to negative biases against AI, reducing credibility, particularly among technology-skeptical audiences (Wang and Huang, 2024; Wang and Ophir, 2024).
@@
AI integration in journalism should be conducted responsibly (Cools and Diakopoulos, 2024). Challenges include addressing algorithmic bias, preventing AI-driven misinformation or “hallucinations”, and safeguarding journalistic autonomy and core values (Cools and Diakopoulos, 2024). News organizations are encouraged to create structures to experiment with, assess, and refine best practices for incorporating AI tools across all stages of news production, from information gathering to editing and distribution (Araujo et al., 2023; Cools and Diakopoulos, 2024). A collaborative model where AI supports rather than replaces journalists, in conjunction with more effective frameworks for sharing information on AI across every level of news organizations, is generally considered best practice (Dodds et al., 2024).
@@ 3 Research Questions and Hypotheses
Our study addresses the following research questions:
@@
How do respondents rate the quality (expertise, readability, and credibility) of AI-assisted and AI-generated news articles, compared to human-generated news articles?
@@
What impact does disclosing AI’s role in generating an article have on respondents’ willingness to read AI-assisted or AI-generated news after they have rated the article for quality?
By answering these two research questions, we extend the literature in three important ways. First, we combine a detailed measurement of “quality” with three realistic levels of AI-involvement in news generation, reflecting current capabilities and practices. Second, we make sure that respondents’ views on AI are not asked in the abstract, which may reflect inaccurate preconceptions, but are rooted in specific examples we provide. Third, we separate quality evaluations from willingness to read AI-generated news by disclosing the degree of AI involvement after participants evaluated the articles’ quality, which ensures that awareness of AI’s role does not affect quality ratings.
The first set of hypotheses focus on quality assessments of news articles with different degrees of AI involvement. Given that a meta-analysis of experimental studies found that human-written news was generally rated higher for quality (Graefe and Bohlken, 2020), our pre-registered hypotheses state that news articles generated with AI involvement are rated lower for all dimensions of quality. However, we acknowledge that more recent studies may challenge these expectations (Wölker and Powell, 2021).
@@
Our second set of hypotheses is designed to explore people’s attitudes once they learn the true role of AI in the article they have read, after rating them for quality. Given that AI-generated content is often met with distrust, particularly when it involves political topics (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020), we expect that respondents in the AI-assisted and AI-generated groups will show a lower willingness to continue reading their articles after being informed of the true generation process.
@@
Our final hypotheses address attitudes towards AI usage in news articles in general, that is, not regarding the particular articles they have read and rated. Respondents will be asked to what degree they would be willing to read AI-generated news in the future. This approach is more in line with opinion-based studies asking abstract questions. However, in our study we ask this question after respondents have engaged with examples of the kinds of articles we are asking about. There is evidence that greater technological competence, AI familiarity, and AI knowledge increases trust in AI (Novozhilova et al., 2024), which suggests that engagement with AI articles may decrease skepticism towards them. Therefore, we expect that respondents in the AI-assisted and AI-generated display express willingness to read AI-generated news articles in the future.
@@ 4 Experimental Design
After informed consent, participants first completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation. Participants were then randomly assigned to one of three conditions: a control group that read excerpts from human-generated articles (“human-generated” group), a first treatment group that read excerpts that were rewritten with the help of ChatGPT (“AI-assisted” group), and another treatment group that read excerpts that were entirely generated by ChatGPT (“AI-generated” group).
To ensure consistency and create a set of texts that is comparable across the three groups, all articles are derived from the human-generated articles that are shown to the control group. We used the following procedure to generate the articles. For AI-assisted articles, we copy-pasted the original article into ChatGPT and asked it to rewrite the article without losing any information. This resulted in articles for the AI-assisted group that are similar to the originals, but are rewritten by the AI system. For AI-generated articles, we only provided ChatGPT with the title and lead of the original article. We then asked ChatGPT to generate a short article in the same style as the source of the human-generated articles.
Each participant read two excerpts, randomly drawn from a pool of ten articles on Swiss politics, with each excerpt having a hard cutoff at 150 words to maintain consistency, avoid bias due to article length, and simulate a paywall for the purposes of our third outcome (willingness to “continue reading” the article). The random selection of two articles from a pool of ten texts for each group minimizes the risk that the outcomes depend on specific topics. Moreover, for each article, we asked respondents to answer a simple question on the article’s content. This allows us to check that respondents have read the text in sufficient detail.
@@
Following the initial rating, participants were then informed about the true creation process of the articles. After learning how the texts they read were actually created, respondents were asked if they would be willing to continue reading the full articles after reviewing the excerpts again. In addition, all respondents were asked to report if they would be willing to read AI-generated news articles in the future, on a 1-5 Likert-scale.

### v3

# Willingness to Read AI-Generated News Is Not Driven by Their Perceived Quality

@@ Abstract
The advancement of artificial intelligence has led to its application in many areas, including news media, which makes it crucial to understand public reception of AI-generated news. This preregistered study investigates (i) the perceived quality of AI-assisted and AI-generated versus human-generated news articles, (ii) whether disclosure of AI’s involvement in generating these news articles influences engagement with them, and (iii) whether such awareness affects the willingness to read AI-generated articles in the future. We conducted a survey experiment with 599 Swiss participants, who evaluated the credibility, readability, and expertise of news articles either written by journalists (control group), rewritten by AI (AI-assisted group), or entirely written by AI (AI-generated group). Our results indicate that all articles were perceived to be of equal quality. When participants in the treatment groups were subsequently made aware of AI’s role, they expressed a higher willingness to continue reading the articles than participants in the control group. However, they were not more willing to read AI-generated news in the future. These results suggest that aversion to AI usage in news media is not primarily rooted in a perceived lack of quality, and that by disclosing using AI, journalists could induce more short-term engagement.
@@ 1 Introduction
The news media landscape is evolving significantly with the advancement of artificial intelligence (AI). AI applications, especially in natural language generation, have become integral to the production of news (Cools and Diakopoulos, 2024; Simon, 2024b, a). The integration of AI in the news presents both opportunities and risks for democracy (Arguedas and Simon, 2024). On the one hand, AI has the potential to make news more accessible and tailored to diverse audiences, supporting informed public debate. On the other hand, however, challenges arise around transparency, bias, and trust, as AI models often reflect the dominant worldviews of their creators, potentially narrowing the range of perspectives in public discourse. Previous research has found that audiences tend to distrust AI-generated content, particularly in politically sensitive contexts, and often prefer human-authored news, even when AI-generated content is evaluated as equally credible (Waddell, 2018; Toff and Simon, 2024; Jia et al., 2024).
@@
A specific issue is the public’s perception of AI-produced news. Understanding attitudes towards AI-generated news is essential as it may directly impact the public’s political knowledge and trust. Indeed, if the public distrusts AI-generated content and chooses not to engage with it, they may miss out on important political information, leading to a less informed citizenry. A recent study revealed that only 29% of Swiss respondents would read fully AI-generated news, 55% would read news if AI assisted in its creation, and 84% would prefer news created without AI involvement (Vogler et al., 2023). This highlights skepticism towards AI-generated news amongst the Swiss population, consistent with surveys from other countries (Fletcher and Nielsen, 2024). Such aversion could be problematic because it could undermine the credibility and effectiveness of news organizations. While previous research has often suggested that audience reluctance stems from concerns about AI-generated news being of lower quality (Waddell, 2018), this assumption has not been directly tested.
@@
In this pre-registered study, conducted with Swiss participants, we investigate the perceived quality of AI-assisted and AI-generated versus human-generated news articles prior to disclosing whether the articles were created by humans, AI-assisted processes, or entirely by AI, and whether subsequent disclosure of the source influences self-reported willingness to engage with those articles as well as the willingness to read AI-assisted or AI-generated articles in the future. Our results indicate that all kinds of news articles, regardless of whether they were written by journalists or (assisted by) AI, were perceived to be of similar quality in terms of credibility, readability, and expertise. This challenges the assumption that skepticism towards AI-generated journalism is driven by concerns about content quality. Instead, our findings align with prior work suggesting that audience resistance is rooted in deeper trust and transparency-related concerns (Barrolleta and Sandoval-Martín, 2024; Bien-Aimé et al., 2025; Jia et al., 2024).
Moreover, while participants who learned about AI’s involvement showed a higher willingness to continue reading the specific article they were presented with—suggesting an initial engagement effect—this did not translate into a greater willingness to consume AI-generated news in the future. This reinforces findings that transparency about AI use can enhance immediate engagement, possibly due to novelty or curiosity, but does not necessarily increase long-term acceptance (Toff and Simon, 2024). These results indicate that public aversion to AI-generated news may not be a matter of perceived content quality but rather a broader issue of trust and perception, highlighting the role of audience heuristics in shaping attitudes towards AI in journalism.
These findings have important implications for both journalism practice and future research. If AI-generated news is perceived to be of equal quality but still met with resistance, efforts to increase public acceptance may need to focus on improving trust rather than simply enhancing AI’s capabilities. Future work should explore the mechanisms behind this resistance, examining how AI transparency, human oversight, and audience familiarity with AI influence perceptions of credibility. Additionally, research should investigate long-term trends in engagement and trust to better understand whether exposure to AI-generated news over time leads to greater acceptance or sustained skepticism. Understanding these dynamics will be crucial for integrating AI into journalism in a way that complements rather than undermines public trust in news media.
2 Perceptions and Challenges of AI-Generated News
@@
AI is increasingly integrated into news production workflows, handling tasks such as data analysis, trend identification, content aggregation, and personalization (Cools and Koliska, 2024). Structured data-based formats like sports and financial reporting are particularly amenable to AI-generated content due to their formulaic nature. AI can assist with data-heavy tasks, enabling journalists to focus on responsibilities that demand human judgment, creativity, and critical thinking, such as investigative reporting and nuanced writing (Clerwall, 2014; Haim and Graefe, 2017). However, concerns persist regarding the opacity of AI-driven processes. Research on The Washington Post’s use of AI-driven news automation shows that while external transparency efforts exist, internal newsroom transparency remains limited, which can lead to errors and hinder journalists’ ability to critically assess AI-generated content (Cools and Koliska, 2024). These applications suggest potential for further AI expansion into other news formats, which may amplify the news industry’s dependency on technological platforms.
Research on public perceptions of news authored by AI reveals a mixed picture. While earlier studies demonstrated preferences for human-written news (Graefe et al., 2018; Graefe and Bohlken, 2020), more recent work found few notable differences in perceived quality, expertise, credibility, and readability between human-written and AI-generated content (Wölker and Powell, 2021; Kim et al., 2020; Kim, Xu and Merrill Jr, 2022). A Turing test-based study assessing Spanish-language AI-generated news similarly found that AI-authored articles were perceived as equally credible to those written by professional journalists, though human authors were still seen as producing more engaging narratives (Barrolleta and Sandoval-Martín, 2024). Nonetheless, AI-produced content tends to be distrusted, especially for politically or emotionally sensitive subjects (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020). Experimental evidence indicates that when news is explicitly labeled as AI-generated, it is often judged as less trustworthy, even when the content itself is evaluated as accurate or fair (Toff and Simon, 2024; Jia et al., 2024). This effect is particularly pronounced among audiences who already hold high levels of trust in news, suggesting that those most engaged with journalism are also most resistant to AI involvement (Toff and Simon, 2024).
Attribution and transparency play a crucial role in shaping audience perceptions of AI-generated news. Research on how readers interpret AI involvement in news articles shows that many assume an exaggerated degree of automation, even when AI plays only a minor role (Bien-Aimé et al., 2025). Furthermore, the way AI authorship is framed in bylines significantly influences credibility assessments; studies have found that the more AI appears to contribute to an article, the lower the perceived message and source credibility (Jia et al., 2024). However, transparency cues such as source disclosure can mitigate some of these negative effects, especially when AI-generated content is accompanied by explicit references to human oversight or lists of data sources used in its creation (Toff and Simon, 2024).
The issue of bias in AI-generated news remains a persistent concern. While some studies suggest AI is perceived as more objective than human journalists (Wölker and Powell, 2021; Wang and Ophir, 2024), this perception may be misleading. AI systems trained on biased data can inadvertently perpetuate and amplify existing biases. Experiments have found that audiences tend to evaluate AI-generated news as less biased, but this effect diminishes when AI involvement is explicitly disclosed, suggesting that perceived neutrality is contingent on a lack of transparency rather than an inherent feature of AI-generated content (Waddell, 2019, 2018).
@@
While transparency is generally seen as beneficial for accountability and audience trust, simply disclosing AI’s involvement may be insufficient or even counterproductive (Wang and Huang, 2024; Wang and Ophir, 2024). People tend to be skeptical of headlines labeled as AI-generated, regardless of whether they are genuinely AI-produced or actually created by humans, partly because they tend to assume an extreme, possibly unrealistic degree of AI involvement (Altay and Gilardi, 2024). Research has also shown that AI-generated content is often seen as more credible when audiences are unaware of its origin, but once labeled as AI-generated, perceptions shift negatively (Jia et al., 2024; Waddell, 2018). This phenomenon suggests that audience heuristics about AI play a crucial role in shaping credibility assessments.
@@
AI integration in journalism should be conducted responsibly (Cools and Koliska, 2024). Challenges include addressing algorithmic bias, preventing AI-driven misinformation or “hallucinations,” and safeguarding journalistic autonomy and core values. Recent findings highlight the need for clearer internal guidelines on AI use within newsrooms, as many organizations still lack standardized attribution policies (Bien-Aimé et al., 2025). News organizations are encouraged to create structures to experiment with, assess, and refine best practices for incorporating AI tools across all stages of news production, from information gathering to editing and distribution (Araujo et al., 2023; Cools and Diakopoulos, 2024). A collaborative model where AI supports rather than replaces journalists, in conjunction with more effective frameworks for sharing information on AI across every level of news organizations, is generally considered best practice (Dodds et al., 2024).
@@ 3 Research Questions and Hypotheses
Building on the existing literature and the gaps identified in previous studies, our research aims to systematically examine how AI involvement in news generation affects audience perceptions. Specifically, we address the following research questions:
@@
Prior to disclosing AI’s role, how do respondents evaluate the quality of AI-assisted and AI-generated news articles in terms of expertise, readability, and credibility, compared to human-generated news articles?
@@
How does disclosing AI’s role in generating an article influence respondents’ willingness to engage with AI-assisted or AI-generated news after they have rated the article’s quality?
These questions allow us to extend prior research in three ways. First, we operationalize “quality” as a multidimensional construct, assessing expertise, readability, and credibility across three distinct levels of AI involvement, reflecting real-world journalistic practices. Second, instead of relying on abstract survey questions about AI, which may be influenced by preconceptions, we ground participants’ responses in concrete examples of AI-generated and AI-assisted news. Third, we separate the evaluation of news article quality from attitudes toward AI-generated news by only revealing AI’s involvement extitafter respondents have rated the articles. This design ensures that initial quality assessments are not biased by prior knowledge of AI authorship.
3.1 Quality Perceptions of AI-Generated News
The first set of hypotheses investigates whether AI-assisted or AI-generated news is perceived as lower in quality compared to human-written articles. Prior studies have found that human-written news is generally rated higher for quality (Graefe and Bohlken, 2020). This aligns with the authority heuristic and social presence heuristic, which suggest that content created by humans is typically seen as more credible due to its perceived human expertise and social presence (Sundar, 2008). Given these established biases, we expect that articles with AI involvement will receive lower quality ratings, though more recent findings challenge this assumption (Wölker and Powell, 2021; Barrolleta and Sandoval-Martín, 2024).
@@ 3.1 Quality Perceptions of AI-Generated News
3.2 Impact of AI Disclosure on Engagement
The second set of hypotheses examines how awareness of AI involvement affects engagement with the article. Previous research suggests that AI-generated content is often met with skepticism, particularly when it covers political topics (Wang and Huang, 2024; Wang and Ophir, 2024; Vogler et al., 2023; Tandoc Jr, Yao and Wu, 2020). We therefore hypothesize that respondents will be less inclined to continue reading once they learn about AI’s role. However, recent findings that disclosure may also spark curiosity or novelty-driven engagement may challenge this premise (Toff and Simon, 2024).
@@ 3.2 Impact of AI Disclosure on Engagement
3.3 Long-Term Attitudes Toward AI-Generated News
The final set of hypotheses assesses how engagement with AI-generated content influences respondents’ general willingness to read AI-generated news in the future. Unlike the previous hypotheses, which examine reactions to specific articles, these hypotheses explore broader attitudes towards AI-generated journalism. While prior surveys indicate high levels of skepticism towards AI-generated news (Vogler et al., 2023), experimental research suggests that direct exposure to AI-generated content may mitigate these concerns, as familiarity with AI often reduces skepticism and increases trust (Novozhilova et al., 2024).
@@ 3.3 Long-Term Attitudes Toward AI-Generated News
By disentangling quality assessments from short-term engagement dynamics and long-term attitudes, our study contributes to a more detailed understanding of how AI-generated journalism is perceived.
@@ 4 Experimental Design
Figure 1: Survey experiment flow
Figure 1 depicts the survey experiment flow: After informed consent, participants first completed a pre-survey to determine their socio-demographic characteristics, including age, gender, education level, and political orientation. Participants were then randomly assigned to one of three conditions: a control group that read excerpts from human-generated articles (“human-generated” group), a first treatment group that read excerpts that were rewritten with the help of ChatGPT (“AI-assisted” group), and another treatment group that read excerpts that were entirely generated by ChatGPT (“AI-generated” group).
To ensure consistency and create a set of texts that is comparable across the three groups, all articles are derived from actual articles on Swiss politics published on the news website of Switzerland’s public broadcaster (SRF), which ensures homogeneity in style, quality, as well as topic to a certain extent (that is, only Swiss politics, with variation in sub-topics). These articles are shown to the control group. We used the following procedure to generate articles with AI involvement. For AI-assisted articles, we copy-pasted the original article into ChatGPT and asked it to rewrite the article without losing any information. This resulted in articles for the AI-assisted group that are similar to the originals, but are rewritten by the AI system. For AI-generated articles, we only provided ChatGPT with the title and lead of the original article. We then asked ChatGPT to generate a short article in the same style as the source of the human-generated articles.
Each participant read two excerpts, randomly drawn from a pool of ten articles for each category (human-generated, AI-assisted, AI-generated), making a total pool of 30 articles across all experimental conditions. Each excerpt had a hard cutoff at 150 words to maintain consistency, avoid bias due to article length, and simulate a paywall for the purposes of our third outcome (willingness to “continue reading” the article). The random selection of two articles from a pool of ten texts for each group minimizes the risk that the outcomes depend on specific topics. Moreover, for each article, we asked respondents to answer a simple question on the article’s content. This allows us to check that respondents have read the text in sufficient detail.
@@
In a second step — following the initial rating — participants were informed about the actual creation process of the articles. After learning how the texts they read were created, respondents were asked if they would be willing to continue reading the full articles after reviewing the excerpts again. In addition, all respondents were asked to report if they would be willing to read AI-generated news articles in the future, on a 1-5 Likert-scale.